# Expected Information Gain Estimation via Density Approximations: Sample Allocation and Dimension Reduction

### v1

@@ 2.2.2 Parameterization and the choice of basis functions
where $g:\mathbb{R}\to\mathbb{R}_{>0}$ is strictly positive and bijective, and satisfies further regularity conditions detailed in Baptista et al. [3]. We substitute this representation of each $S^{k}$ into (14) and solve the resulting unconstrained minimization problem, $\min_{f}\widehat{J}_{k}(\mathcal{R}_{k}(f))$ over some linear space of functions $\mathcal{V}_{k}\ni f$. As explained in Baptista et al. [3], this construction provides important optimization guarantees: there are no spurious local minima (every local minimizer $f^{\ast}$ is global minimizer) and under certain tail conditions on $\pi$ and $f$, we provably recover the true KR rearrangement.
Here, we parameterize $f$ using a linear expansion of tensor-product Hermite polynomials $\psi_{\alpha}$ as $f(z_{1:k})=\sum_{\alpha}c_{\alpha}\psi_{\alpha}(z_{1:k})$ and solve for its coefficients $c_{\alpha}$. We refine this polynomial space using a greedy adaptive procedure and apply cross-validation to select the total number of polynomial basis functions to retain. We refer to Baptista et al. [3] for complete details on these numerical procedures.
We note also that while our focus thus far in Section 2.2 and Subsection 2.2.2, and in the numerical experiments reported later in this paper, is on strictly triangular maps, we will see in Section 2.2.3 that block triangular maps as described in Baptista et al. [5] are also suitable for estimating the marginal and conditional densities relevant to the OED problem. In contrast with the strictly triangular case, there is no unique choice of block-triangular map. One natural choice is a conditional Brenier map, as described in Carlier et al. [12], Baptista et al. [5], Wang et al. [66]. Another broad family of choices are given by conditional normalizing flows, which are specific parameterizations of the bottom block of a block-triangular map. We use strictly triangular maps in our numerical experiments because the associated computations are relatively fast, and because they enjoy the optimization guarantees mentioned above; to our knowledge, such guarantees are not available for other continuous map representations.
@@ 2.2.3 From marginal and conditional density estimation to EIG bounds
where $S^{\mathcal{Y}}\colon\mathbb{R}^{n_{y}}\rightarrow\mathbb{R}^{n_{y}}$ and $S^{\mathcal{X}}\colon\mathbb{R}^{n_{y}+n_{x}}\to\mathbb{R}^{n_{x}}$. We choose $\eta$ to be standard Gaussian on $\mathbb{R}^{n_{y}+n_{x}}$. Our parameterizations ensure that $S^{\mathcal{Y}}$ is invertible and that $\nabla_{y}S^{\mathcal{Y}}$ exists, and similarly that $x\mapsto S^{\mathcal{X}}(y,x)$ is invertible and that $\nabla_{x}S^{\mathcal{X}}(y,\cdot)$ exists for every $y\in\mathbb{R}^{n_{y}}$; this applies specifically to the numerical representations of triangular maps mentioned above. Note that such properties are satisfied almost everywhere by the KR map for absolutely continuous distributions with full support; see Spantini et al. [60, Sec. 3] and Santambrogio [57, Sec. 2.3] for a discussion.
If we select our reference distribution $\eta$ to be any distribution whose density factorizes as $\eta(x,y)=\eta_{X}(x)\eta_{Y}(y)$ (which clearly includes the standard normal), Baptista et al. [5, Theorem 2.4] shows that we can express the marginal density of $Y$ as $\pi_{Y}=(S^{\mathcal{Y}})^{\sharp}\eta_{Y}$ and the posterior density of $X$ as $\pi_{X|Y=y}=S^{\mathcal{X}}(y,\cdot)^{\sharp}\eta_{X}$ for any $y\in\mathbb{R}^{n_{y}}$.
@@
$\widehat{\pi}_{X|Y=y}$ $=\widehat{S}^{\mathcal{X}}(y,\cdot)^{\sharp}\eta_{X},$
@@
$\widehat{\pi}_{Y|X=x}$ $=\widehat{U}^{\mathcal{Y}}(x,\cdot)^{\sharp}\eta_{Y}.$
@@ 2.3 Convergence analysis and optimal sample allocation
Given that the transport maps are estimated using $N$ training samples (see (14)), and the empirical average is computed using $M$ evaluation samples (see (6), (7), (8), (9)), we aim to investigate the optimal sample allocation between the number of training samples $N$ and the number of evaluation samples $M$, given a total of $L=N+M$ samples, with the objective of minimizing the mean squared error (MSE) of the EIG estimator. Here, we use $\widehat{\mathrm{EIG}}_{M,N}$ to indicate that the EIG estimator is obtained using $N$ training samples and $M$ evaluation samples. For an EIG estimator $\widehat{\mathrm{EIG}}_{M,N}$, the bias is defined as
@@
We refer readers to Section 10.6.2 of Casella and Berger [13] for the regularity conditions that guarantee the consistency and asymptotic normality of the MLE. We now present our main result on the bias and variance of a resulting EIG estimator constructed using an approximate density based on transport maps.
@@
Under Assumptions 2.3 and 2.3, let $g(\alpha)=\mathbb{E}_{\pi_{Y}}\left[\log q_{Y}(y;\alpha)\right]$. Suppose $g$ is continuous and bounded, and that both $\nabla g(\alpha),\nabla^{2}g(\alpha)$ exist and are continuous. Then, the EIG estimator
@@
has bias converging to zero at the rate of $O(1/N)$ and variance converging to zero at the rate of $O(1/M+1/N^{2})$. Thus, the MSE converges at the rate of $O(1/M+1/N^{2})$.
The proof of Theorem 2.1 is presented in Appendix B.1. As discussed earlier, we can also approximate the posterior distribution, assuming the prior distribution is known. The analysis is similar and the convergence rate of the MSE is also $O(1/M+1/N^{2})$. We can adopt the same approach when it is necessary to approximate both the likelihood and the marginal distribution of $Y$. As in (16), we define $\mathcal{Q}^{\mathcal{Y}|X}$ as
@@
The maximum likelihood estimator $\widehat{\beta}_{N}$ is consistent and asymptotically normal. We then have the following theorem.
@@
Under Assumptions 2.32.32.3 and 2.3, and assuming $q_{Y}\in\mathcal{Q}^{\mathcal{Y}}$ and $q_{Y|X}\in\mathcal{Q}^{\mathcal{Y|X}}$, and let $h(\beta)=\mathbb{E}_{\pi_{X,Y}}\left[\log q_{Y|X}(y|x;\beta)\right]$. Suppose $g$ and $h$ are both continuous and bounded, and that $\nabla g(\alpha),\nabla^{2}g(\alpha)$ and $\nabla h(\beta),\nabla^{2}h(\beta)$ exist and are continuous in some neighborhood of $\alpha^{*}$ and $\beta^{*}$, respectively. Then bias of the EIG estimator
@@
In the context of Theorem  2.1 and Theorem  2.2, given $L$ samples from the joint distribution $\pi_{X,Y}$, the optimal allocation between the training and evaluation samples should be set as $\frac{M}{N}\sim O(L^{1/3})$. Correspondingly, the optimal convergence rate of the MSE is $O(L^{-1})$.
The proof of this corollary is provided in Appendix B.3. We observe that if $\frac{M}{N}\sim O(L^{p})$ for some $p>0$, the MSE decreases at a rate of $2L^{p-2}+L^{-p-1}+L^{2p-2}+L^{-2}+L^{-1}$. If we choose $p>1/2$, the leading term will have an order greater than $-1$. An asymptotic optimal convergence rate of $O(L^{-1})$ is thus achieved for $p\in(0,\frac{1}{2}]$. This is confirmed in the numerical examples presented in the Section 4.1.
The final comment we would like to make in this section is that this result only applies when the map is in-class, and the optimization problem is solved by dropping all constants in the asymptotic notation Therefore, parameter tuning may still be necessary in practice to achieve the optimal sample allocation to minimize the MSE.
@@ 3 EIG estimation in high dimensions
When the dimension of the parameters and data is high, directly estimating MI becomes computationally expensive. To address this, we aim to reduce the dimensionality of the parameters and data simultaneously, ensuring that the lower-dimensional subspace still preserves the mutual information. In this section, we apply the dimension reduction methods proposed in Baptista et al. [2]. Given two unitary matrices $U=[U_{r},U_{\perp}]\in\mathbb{R}^{n_{x}\times n_{x}}$ and $V=[V_{s},V_{\perp}]\in\mathbb{R}^{n_{y}\times n_{y}}$, consider the decompositions
@@
To find rotation matrices $U$ and $V$ and reduced dimensions $r,s$ that minimize the posterior approximation error and satisfy the bound above, we use the following theorem from Baptista et al. [2].
@@
Let $X$ and $Y$ be random variables whose joint distribution satisfies the logarithmic Sobolev inequality with constant $C(\pi_{X,Y})$ as stated by Definition 2 in Baptista et al. [2]. Then, we have
@@ 3.1 Gaussian prior and likelihood case
Let $\widetilde{U}$ and $\widetilde{V}$ denote the matrices whose columns are the eigenvectors of $H_{\widetilde{X}}$ and $H_{\widetilde{Y}}$, respectively. Then, using
@@
we obtain the matrices $U$ and $V$, and $\widetilde{U}$ and $\widetilde{V}$ satisfy the properties $\widetilde{U}^{\top}\Sigma_{X}\widetilde{U}=I$ and $\widetilde{V}^{\top}\Sigma_{\mathcal{E}}\widetilde{V}=I$. Let $U_{r}$ and $V_{s}$ be the matrices composed of the first $r$ and $s$ columns of $U$ and $V$, respectively. The optimal projection, in terms of preserving mutual information, is given by $X_{r}=U_{r}^{\top}X$ and $Y_{s}=V_{s}^{\top}Y$. We can then compute the mutual information $\mathcal{I}(X_{r},Y_{s})$ in the reduced coordinates, and the error is proportional to $\sum_{i=r+1}^{n_{x}}\lambda_{i}\left(U^{\top}_{\perp}H_{X}U_{\perp}\right)+\sum_{i=s+1}^{n_{y}}\lambda_{i}\left(V^{\top}_{\perp}H_{Y}V_{\perp}\right)$, where $\lambda_{i}$ denotes the $i$th largest eigenvalue of the corresponding matrix. For further details, see Baptista et al. [2].
@@ 3.2 Lower bounds of MI and the choice of estimators
To reduce the dimensionality of the samples, we need access to the gradient of the forward model, specifically $\nabla G(X)$, which is not available in a completely likelihood-free setting. This limits us to using $\widehat{\textrm{EIG}}_{\mathrm{m}}$ and $\widehat{\textrm{EIG}}_{\mathrm{pos}}$. In the rotated space, we define:
$\widehat{\textrm{EIG}}_{\mathrm{m}}(r,s)$ $=\frac{1}{M}\sum_{i=1}^{M}\left(\log\frac{\pi_{Y_{s}|X_{r}}(y_{s}^{i}|x_{r}^{i})}{\widehat{\pi}_{Y_{s}}(y_{s}^{i})}\right),$
$\widehat{\textrm{EIG}}_{\mathrm{pos}}(r,s)$ $=\frac{1}{M}\sum_{i=1}^{M}\left(\log\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}|y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}\right).$
Then we have,
$\widehat{\textrm{EIG}}_{\mathrm{pos}}(r,s)\approx\mathbb{E}_{\pi_{X_{r},Y_{s}}}\left[\log\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}|y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}\right]\leq\mathcal{I}(X_{r};Y_{s})\leq\mathcal{I}(X;Y),$
where the first inequality follows from the lower bound of the EIG, and the second is obtained from the data processing inequality. However, the relationship between $\widehat{\textrm{EIG}}_{\mathrm{m}}(r,s)$ and the true EIG remains unclear. Although $\mathbb{E}_{\pi_{X_{r},Y_{s}}}\left[\log\frac{\pi_{Y_{s}|X_{r}}(y_{s}^{i}|x_{r}^{i})}{\widehat{\pi}_{Y_{s}}(y_{s}^{i})}\right]\geq\mathcal{I}(X_{r},Y_{s})$, it does not provide a clear upper or lower bound for $\mathcal{I}(X,Y)$. Therefore, we focus on computing $\widehat{\textrm{EIG}}_{\mathrm{pos}}$ in high-dimensional settings. We summarize the proposed method for EIG estimation using low-dimensional projections in Algorithm 2.
@@
7:   Use $M$ pairs of samples to estimate EIG by evaluating the empirical average $\widehat{\textrm{EIG}}_{\mathrm{pos}}=\frac{1}{M}\sum_{i=1}^{M}\left(\log\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}\mid y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}\right)$.

### v2

@@ 2.2.2 Parameterization and the choice of basis functions
where $g:\mathbb{R}\to\mathbb{R}_{>0}$ is strictly positive and bijective, and satisfies further regularity conditions detailed in [5]. We substitute this representation of each $S^{k}$ into (14) and solve the resulting unconstrained minimization problem, $\min_{f}\widehat{J}_{k}(\mathcal{R}_{k}(f))$ over some linear space of functions $\mathcal{V}_{k}\ni f$. As explained in [5], this construction provides important optimization guarantees: there are no spurious local minima (every local minimizer $f^{\ast}$ is global minimizer) and under certain tail conditions on $\pi$ and $f$, we provably recover the true KR rearrangement.
Here, we parameterize $f$ using a linear expansion of tensor-product Hermite polynomials $\psi_{\alpha}$ as $f(z_{1:k})=\sum_{\alpha}c_{\alpha}\psi_{\alpha}(z_{1:k})$ and solve for its coefficients $c_{\alpha}$. We refine this polynomial space using a greedy adaptive procedure and apply cross-validation to select the total number of polynomial basis functions to retain. We refer to [5] for complete details on these numerical procedures.
We note also that while our focus thus far in Section 2.2 and Subsection 2.2.2, and in the numerical experiments reported later in this paper, is on strictly triangular maps, we will see in Section 2.2.3 that block triangular maps as described in [3] are also suitable for estimating the marginal and conditional densities relevant to the OED problem. In contrast with the strictly triangular case, there is no unique choice of block-triangular map. One natural choice is a conditional Brenier map, as described in [12, 3, 67]. Another broad family of choices are given by conditional normalizing flows, which are specific parameterizations of the bottom block of a block-triangular map. We use strictly triangular maps in our numerical experiments because the associated computations are relatively fast, and because they enjoy the optimization guarantees mentioned above; to our knowledge, such guarantees are not available for other continuous map representations.
@@ 2.2.3 From marginal and conditional density estimation to EIG bounds
where $S^{\mathcal{Y}}\colon\mathbb{R}^{n_{y}}\rightarrow\mathbb{R}^{n_{y}}$ and $S^{\mathcal{X}}\colon\mathbb{R}^{n_{y}+n_{x}}\to\mathbb{R}^{n_{x}}$. We choose $\eta$ to be standard Gaussian on $\mathbb{R}^{n_{y}+n_{x}}$. Our parameterizations ensure that $S^{\mathcal{Y}}$ is invertible and that $\nabla_{y}S^{\mathcal{Y}}$ exists, and similarly that $x\mapsto S^{\mathcal{X}}(y,x)$ is invertible and that $\nabla_{x}S^{\mathcal{X}}(y,x)$ exists for every $y\in\mathbb{R}^{n_{y}}$; this applies specifically to the numerical representations of triangular maps mentioned above. Note that such properties are satisfied almost everywhere by the KR map for absolutely continuous distributions with full support; see [59, Sec. 3] and [57, Sec. 2.3] for a discussion.
If we select our reference distribution $\eta$ to be any distribution whose density factorizes as $\eta(x,y)=\eta_{X}(x)\eta_{Y}(y)$ (which clearly includes the standard normal), [3, Theorem 2.4] shows that we can express the marginal density of $Y$ as $\pi_{Y}=(S^{\mathcal{Y}})^{\sharp}\eta_{Y}$ and the posterior density of $X$ as $\pi_{X|Y=y}=S^{\mathcal{X}}(y,\cdot)^{\sharp}\eta_{X}$ for any $y\in\mathbb{R}^{n_{y}}$.
@@
$\widehat{\pi}_{X|Y=y}$ $=\left(\widehat{S}^{\mathcal{X}}(y,\cdot)\right)^{\sharp}\eta_{X},$
@@
$\widehat{\pi}_{Y|X=x}$ $=\left(\widehat{U}^{\mathcal{Y}}(x,\cdot)\right)^{\sharp}\eta_{Y}.$
@@ 2.3 Convergence analysis and optimal sample allocation
Given that the transport maps are estimated using $N$ training samples (see (14)), and the empirical average is computed using $M$ evaluation samples (see (6), (7), (8), (9)), we aim to investigate the optimal sample allocation between the number of training samples $N$ and the number of evaluation samples $M$, given a total of $L=N+M$ samples, with the objective of minimizing the MSE of the EIG estimator. Here, we use $\widehat{\mathrm{EIG}}_{M,N}$ to indicate that the EIG estimator is obtained using $N$ training samples and $M$ evaluation samples. For an EIG estimator $\widehat{\mathrm{EIG}}_{M,N}$, the bias is defined as
@@
We refer the reader to Section 10.6.2 of [13] for the regularity conditions that guarantee the consistency and asymptotic normality of the MLE. We now present our main result on the bias and variance of an resulting EIG estimator constructed using an approximate density based on transport maps.
@@
Under the Assumptions 2.3 and 2.3, let $g(\alpha)=\mathbb{E}_{\pi_{Y}}\left[\log q_{Y}(y;\alpha)\right]$, where $q_{Y}\in\mathcal{Q}^{\mathcal{Y}}$. Suppose $g$ is continuous and bounded, and that both $\nabla g(\alpha),\nabla^{2}g(\alpha)$ exist and are continuous. Then, the EIG estimator
@@
has bias converging to zero at a rate of $O(1/N)$ and variance converging to zero at a rate of $O(1/M+1/N^{2})$. Thus, the MSE converges at a rate of $O(1/M+1/N^{2})$.
The proof of Theorem 2.1 is presented in Appendix B.1. We then define $\mathcal{Q}^{\mathcal{Y}|X}$ as
@@
The maximum likelihood estimator $\widehat{\beta}_{N}$ is consistent and asymptotically normal.
@@
Under the Assumptions 2.3 and  2.3, let $h(\beta)=\mathbb{E}_{\pi_{X,Y}}\left[\log q_{Y|X}(y|x;\beta)\right]$, where $q_{Y|X}\in\mathcal{Q}^{\mathcal{Y|X}}$. Suppose $h$ is continuous and bounded, and that both $\nabla h(\beta),\nabla^{2}h(\beta)$ exist and are continuous. Then, the EIG estimator $\frac{1}{M}\sum_{i=1}^{M}\log\frac{q_{Y|X}(y^{i}|x^{i};\widehat{\beta}_{N})}{\pi_{Y}(y^{i})}$ has bias converging to zero at a rate of $O(1/N)$ and variance converging to zero at a rate of $O(1/M+1/N^{2})$. Thus, the MSE converges at a rate of $O(1/M+1/N^{2})$.
Theorem 2.3.
Under the Assumptions 2.32.32.3 and 2.3, suppose $g$ and $h$ are both continuous and bounded, and that $\nabla g(\alpha),\nabla^{2}g(\alpha)$ and $\nabla h(\beta),\nabla^{2}h(\beta)$ exist and are continuous in some neighborhood of $\alpha^{*}$ and $\beta^{*}$, respectively. Then bias of the EIG estimator
@@
In the context of Theorem  2.1 and Theorem  2.3, given $L$ samples from the joint distribution $\pi_{X,Y}$, the optimal allocation between the training and evaluation samples should be set as $\frac{M}{N}\sim O(L^{1/3})$. Correspondingly, the optimal convergence rate of the MSE is $O(L^{-1})$.
The proof of this corollary is provided in Appendix B.4. We observe that if $\frac{M}{N}\sim O(L^{p})$ for some $p>0$, the mean squared error (MSE) decreases at a rate of $2L^{p-2}+L^{-p-1}+L^{2p-2}+L^{-2}+L^{-1}$. If $p>1/2$, the leading term has an order greater than $-1$. For all $p\in(0,\frac{1}{2}]$, the asymptotic convergence rate remains $O(L^{-1})$. The optimal value, $p=1/3$, is obtained by solving a root-finding problem (see Appendix B.4 for more detail), where the difference lies in the order of the non-leading term. The rate of convergence is confirmed in the numerical examples presented in Section 4.1.
@@ 3 EIG estimation in high dimensions
When the dimension of the parameters and data is high, directly estimating MI becomes computationally expensive. To address this, we aim to reduce the dimensionality of the parameters and data simultaneously, ensuring that the lower-dimensional subspace still preserves the mutual information. In this section, we apply the dimension reduction methods proposed in [4]. Given two unitary matrices $U=[U_{r},U_{\perp}]\in\mathbb{R}^{n_{x}\times n_{x}}$ and $V=[V_{s},V_{\perp}]\in\mathbb{R}^{n_{y}\times n_{y}}$, consider the decompositions
@@
To find rotation matrices $U$ and $V$ and reduced dimensions $r,s$ that minimize the posterior approximation error and satisfy the bound above, we use the following theorem from [4].
@@
Let $X$ and $Y$ be random variables whose joint distribution satisfies the logarithmic Sobolev inequality with constant $C(\pi_{X,Y})$ as stated by Definition 2 in [4]. Then, we have
@@ 3.1 Gaussian prior and likelihood case
Let $\widetilde{U}$ and $\widetilde{V}$ denote the matrices whose columns are the eigenvectors of $H_{\widetilde{X}}$ and $H_{\widetilde{Y}}$, respectively. Then, we obtain the matrices $U$ and $V$ using
@@
where $\widetilde{U}$ and $\widetilde{V}$ satisfy the properties $\widetilde{U}^{\top}\Sigma_{X}\widetilde{U}=I$ and $\widetilde{V}^{\top}\Sigma_{\mathcal{E}}\widetilde{V}=I$. Let $U_{r}$ and $V_{s}$ be the matrices composed of the first $r$ and $s$ columns of $U$ and $V$, respectively. The optimal projection, in terms of preserving mutual information, is given by $X_{r}=U_{r}^{\top}X$ and $Y_{s}=V_{s}^{\top}Y$. We can then compute the mutual information $\mathcal{I}(X_{r},Y_{s})$ in the reduced coordinates, and the error is proportional to $\sum_{i=r+1}^{n_{x}}\lambda_{i}\left(U^{\top}_{\perp}H_{X}U_{\perp}\right)+\sum_{i=s+1}^{n_{y}}\lambda_{i}\left(V^{\top}_{\perp}H_{Y}V_{\perp}\right)$, where $\lambda_{i}$ denotes the $i$th largest eigenvalue of the corresponding matrix. For further details, see [4].
@@ 3.2 Lower bounds of MI and the choice of estimators
To reduce the dimensionality of the samples, we need access to the gradient of the forward model, specifically $\nabla G(X)$, which is not available in a completely likelihood-free setting. This limits us to using $\widehat{\mathrm{EIG}}_{\mathrm{m}}$ and $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$. In the rotated space, we define:
$\widehat{\mathrm{EIG}}_{\mathrm{m}}(r,s)$ $=\frac{1}{M}\sum_{i=1}^{M}\log\frac{\pi_{Y_{s}|X_{r}}(y_{s}^{i}|x_{r}^{i})}{\widehat{\pi}_{Y_{s}}(y_{s}^{i})},$
$\widehat{\textrm{EIG}}_{\mathrm{pos}}(r,s)$ $=\frac{1}{M}\sum_{i=1}^{M}\log\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}|y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}.$
Then note that $\mathbb{E}_{\pi_{X_{r},Y_{s}}}\left[\log\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}|y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}\right]\leq\mathcal{I}(X_{r};Y_{s})\leq\mathcal{I}(X;Y)$, where the first inequality follows from the lower bound of the EIG, and the second is obtained from the data processing inequality. Therefore, $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$ serves as a lower bound in the asymptotic sense. However, the relationship between $\widehat{\mathrm{EIG}}_{\mathrm{m}}(r,s)$ and the true EIG remains unclear. Although $\mathbb{E}_{\pi_{X_{r},Y_{s}}}\left[\log\frac{\pi_{Y_{s}|X_{r}}(y_{s}^{i}|x_{r}^{i})}{\widehat{\pi}_{Y_{s}}(y_{s}^{i})}\right]\geq\mathcal{I}(X_{r},Y_{s})$, it does not provide a clear upper or lower bound for $\mathcal{I}(X,Y)$. We thus focus on computing $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$ in high-dimensional settings. We summarize the proposed method for EIG estimation using low-dimensional projections in Algorithm 2.
@@
7:   Use $M$ pairs of samples to estimate EIG by evaluating the empirical average $\widehat{\textrm{EIG}}_{\mathrm{pos}}=\frac{1}{M}\sum_{i=1}^{M}\log\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}\mid y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}$.